# OscNet v1.5: Energy Efficient Hopfield Network on CMOS Oscillators for Image Classification

### v1

@@ Abstract
Machine learning has achieved remarkable advancements but at the cost of significant computational resources. This has created an urgent need for a novel and energy-efficient computational fabric. CMOS Oscillator Networks (OscNet) is a brain inspired and specially designed hardware for low energy consumption. In this paper, we propose a Hopfield Network based machine learning algorithm that can be implemented on OscNet. The network is trained using forward propagation alone to learn sparsely connected weights, yet achieves an $8\%$ improvement in accuracy compared to conventional deep learning models on MNIST dataset. OscNet v1.5 achieves competitive accuracy on MNIST and is well-suited for implementation using CMOS-compatible ring oscillator arrays with Subharmonic Injection Locking (SHIL). In oscillator-based implementation, we utilize only $24\%$ of the connections used in a fully connected Hopfield network, with merely a $0.1\%$ drop in accuracy. OscNet v1.5 relies solely on forward propagation and employs sparse connections, making it an energy-efficient machine learning pipeline designed for CMOS oscillator computing. The repository for OscNet family is: https://github.com/RussRobin/OscNet.
@@ 1 Introduction
Gradient descent based machine learning has achieved significant success in various fields [6, 12, 14]. However, the energy consumption of current von Neumann architecture based CPU and GPU computing is becoming increasingly high and unsustainable [19, 4, 9, 11]. In contrast, the human brain can perform complex tasks under extremely low power, even though the efficiency of neural signal transmission is far lower than that of electric currents in CMOS circuits. Inspired by this, the Complementary Metal-Oxide-Semiconductor (CMOS) Oscillator Network (OscNet) was developed, where information is directly stored in the phase of each oscillator, and inter-oscillator currents model communication between neurons. This architecture is especially efficient on forward propagation in machine learning. Prior work has designed Hebbian learning schemes [7] and specific algorithms for solving nondeterministic polynomial-time (NP) problems [24] on OscNet. Building on this foundation, we propose OscNet v1.5, a brain-inspired and energy-efficient Hopfield Network–style algorithm to support learning-based generalization from original Hopfield Network to image classification tasks [3, 1].
A simplified illustration of the proposed algorithm is shown in Fig. 1. We develop a biologically plausible pipeline for Hopfield-based learning and prediction that relies solely on local information and online learning, consistent with the constraints observed in real neural systems. Training of OscNet is based on forward propagation and Hebbian learning only [2, 13, 7], so it is more energy efficient then current deep learning networks which rely heavily on forward and backward propagation.
Traditional Hopfield Networks are fully connected, which does not reflect the connectivity structure observed in the human brain. In this paper, we construct sparsely connected Hopfield Networks. Sparsely connected architectures can outperform fully connected ones in classification tasks—in terms of both accuracy and computational efficiency—while also being more biologically plausible. We propose a Hebbian style learning rule that determines not only the connection weights but also whether a connection should exist between nodes. Implementing the Hopfield Network on oscillators, we use only $24\%$ of connections while the performance only drops by $0.1\%$ compared to a fully connected Hopfield Network. Our algorithm is energy efficient because it learns from forward propagation only, without the need to do back propagation or gradient descent. Additionally, it is a sparsely connected network, which further conserves energy. The main contributions of this paper are:
  • We proposed OscNet v1.5, a machine learning pipeline designed for CMOS oscillator computing. OscNet v1.5 is energy efficient as it relies solely on forward propagation and employs sparse connections.
@@
  • In oscillator-based implementation, we utilize only $24\%$ of the connections used in a fully connected Hopfield network, with merely a $0.1\%$ drop in accuracy.
@@ 2.2 CMOS Oscillator Networks
CMOS Oscillator Network (OscNet) [7] operates with high-order injection locking, with local connections to neighboring oscillators. Each oscillator is injected with a master pump signal at $N\textperiodcentered f_{0}$, where $f_{0}$ is oscillators’ natural oscillator frequency, and $N$ is a chosen integral. The phases of oscillators can be $n\frac{2\pi}{N}$, where $n=1,2...N-1$. This network mimics the activity of polychronous spiking neurons in the brain [16], where each oscillator represents a single brain cell. In OscNet, all oscillators oscillate at the same frequency, and the information is encoded in the phases of the individual oscillators. The voltage across the oscillator $i$ at time $t$ is:
@@ 3.3 Normalization
Figure 3: Up: examples of training images of class 0,1,2,3 in MNIST [17]. Down: learned features of pixel (0,0) in Hopfield Networks for class 0,1,2,3. The colormap shows absolute value of similarity matrix. Black: pixel (0,0), green: pixel (14,14), and blue: pixel (0,14). The similarity of (0,0)-(0,14), as shown with blue line, is always large across labels, which brings little information in image classification. (0,0)-(0,14) connection in green tells between class 0 and 1. The weights shown are not normalized.

### v2

# OscNet v1.5: Hopfield Network Designed for Image Classification by Energy-Efficient Oscillator Fabrics

@@ Abstract
Machine learning has achieved remarkable advancements but at the cost of significant computational resources. This has created an urgent need for a novel and energy-efficient computational fabric and corresponding algorithms. CMOS Oscillator Networks (OscNet) is a brain inspired and specially designed hardware for low energy consumption. In this paper, we propose a Hopfield Network based machine learning algorithm that can be implemented on OscNet. The network is trained using forward propagation alone to learn sparsely connected weights, yet achieves an $8\%$ improvement in accuracy compared to conventional deep learning models on MNIST dataset. OscNet v1.5 achieves competitive accuracy on MNIST and is well-suited for implementation using CMOS-compatible ring oscillator arrays with Subharmonic Injection Locking (SHIL). In oscillator-based inference, we utilize only $24\%$ of the connections used in a fully connected Hopfield network, with merely a $0.1\%$ drop in accuracy. OscNet v1.5 relies solely on forward propagation and employs sparse connections, making it an energy-efficient machine learning pipeline designed for oscillator computing fabric. The repository for OscNet family is: https://github.com/RussRobin/OscNet.
@@ 1 Introduction
Gradient descent based machine learning has achieved significant success in various fields [1, 2, 3]. However, the energy consumption of current von Neumann architecture based CPU and GPU computing is becoming increasingly high and unsustainable [4, 5, 6, 7]. In contrast, the human brain can perform complex tasks under extremely low power, even though the efficiency of neural signal transmission is far lower than that of electric currents in CMOS circuits. Inspired by this, the Complementary Metal-Oxide-Semiconductor (CMOS) Oscillator Network (OscNet) was developed, where information is directly stored in the phase of each oscillator, and inter-oscillator currents model communication between neurons. This architecture is especially efficient on forward propagation in machine learning. Prior work [8] developed a mathematical model to optimize graph coloring using oscillator circuits, and [9] designed Hebbian learning schemes on OscNet. Building on this foundation, we propose OscNet v1.5, a brain-inspired and energy-efficient Hopfield Network–style algorithm to support learning-based generalization from original Hopfield Network to image classification tasks [10, 11], designed for the oscillator computing fabric.
A simplified illustration of the proposed algorithm is shown in Fig. 1. We develop a biologically plausible pipeline for Hopfield-based learning and prediction that relies solely on local information and online learning, consistent with the constraints observed in real neural systems. Training of OscNet v1.5 is based on forward propagation only [12, 13, 9], so it is more energy efficient then current deep learning networks which rely heavily on forward and backward propagation.
Traditional Hopfield Networks are fully connected, which does not reflect the connectivity structure observed in the human brain. In this paper, we construct sparsely connected Hopfield Networks. Our sparsely connected architecture can outperform fully connected ones in classification tasks—in terms of both accuracy and computational efficiency—while also being more biologically plausible. We propose a Hebbian style learning rule that determines not only the connection weights but also whether a connection should exist between nodes. In oscillator based inference, we use only $24\%$ of connections while the performance only drops by $0.1\%$ compared to a fully connected Hopfield Network. Our algorithm is energy efficient because it learns from forward propagation only, without the need to do back propagation or gradient descent. Additionally, it is a sparsely connected network, which further conserves energy. The main contributions of this paper are:
  • We proposed OscNet v1.5, a machine learning pipeline designed for oscillator computing fabric. OscNet v1.5 is energy efficient as it relies solely on forward propagation and employs sparse connections.
@@
  • In oscillator-based inference, we utilize only $24\%$ of the connections used in a fully connected Hopfield network, with merely a $0.1\%$ drop in accuracy.
@@ 2.2 CMOS Oscillator Networks
In this section, we review the mathematical theory established in [8], upon which OscNet [9] builds. CMOS Oscillator Network (OscNet) [9] operates with high-order injection locking, with local connections to neighboring oscillators. Each oscillator is injected with a master pump signal at $N\textperiodcentered f_{0}$, where $f_{0}$ is oscillators’ natural oscillator frequency, and $N$ is a chosen integral. The phases of oscillators can be $n\frac{2\pi}{N}$, where $n=1,2...N-1$. This network mimics the activity of polychronous spiking neurons in the brain [17], where each oscillator represents a single brain cell. In OscNet, all oscillators oscillate at the same frequency, and the information is encoded in the phases of the individual oscillators. The voltage across the oscillator $i$ at time $t$ is:
@@ 3.3 Normalization
Figure 3: Up: examples of training images of class 0,1,2,3 in MNIST [14]. Down: learned features of pixel (0,0) in Hopfield Networks for class 0,1,2,3. The colormap shows absolute value of similarity matrix. Black: pixel (0,0), green: pixel (14,14), and blue: pixel (0,14). The similarity of (0,0)-(0,14), as shown with blue line, is always large across labels, which brings little information in image classification. (0,0)-(14,14) connection in green tells between class 0 and 1. The weights shown are not normalized.